Find the intervals on which the function f defined by f( x )=x⁴-8x²+16is concave upward or concave downward.
1
Expert's answer
2020-03-12T12:44:18-0400

"f'=4x^3-16x"

"f''=12x^2-16"

"12x^2-16>0\iff 3x^2-4>0\implies x\in (-\infty,-\frac{2\sqrt3 }{3})\bigcup(\frac{2\sqrt3 }{3},+\infty)", then the function is concave upward.

"f''<0\iff x\in (-\frac{2\sqrt3}{3},\frac{2\sqrt3}{3})" , then the function is concave downward.